# Teaching LLMs According to Their Aptitude: Adaptive Reasoning for Mathematical Problem Solving

### v2

@@ 4.1 Experimental Setup
We include two widely used SFT approaches as our baselines: SFT exclusively on CoT data and SFT exclusively on TIR data. We also investigate several selection methods in Section 4.3.
@@ 4.2 Main Results
Results presented in Table 1 demonstrate the effectiveness of our proposed TATA framework, which combines the strengths of CoT and TIR methods for mathematical reasoning tasks. Across various base models, model sizes, and benchmarks, TATA consistently achieves competitive or superior performance compared to standalone CoT and TIR approaches, highlighting its ability to leverage the complementary advantages of both methods.
In-Domain Performance
For in-domain tasks (GSM8K and MATH), TATA achieves the highest average accuracy (ID AVG) in most cases, outperforming both CoT and TIR individually. This suggests that the integration of CoT’s step-by-step reasoning with TIR’s tool-assisted computations enhances the model’s ability to solve problems within its training domain. Notably, for larger models such as Qwen2.5-7B and Qwen2.5Math-7B, TATA achieves ID AVG scores of 78.2% and 81.4%, respectively, demonstrating its scalability and effectiveness as model capacity increases.
Out-of-Domain (OOD) Performance
The OOD results further underscore the robustness of TATA. Across diverse benchmarks such as MAWPS, SVAMP, CollegeMath, and OlympiadBench, TATA consistently achieves the highest OOD average scores, indicating its generalizability.
@@
The results in Table 2 demonstrate that our TATA not only improves accuracy but also enhances inference efficiency compared to standalone CoT and TIR methods. More detailed results are presented in Table 10. Across all model sizes, TATA achieves higher accuracy while maintaining lower token usage and fewer code executions than TIR, and it significantly reduces computational overhead compared to TIR without sacrificing the benefits of tool integration. For instance, with Qwen2.5-7B, TATA achieves a 2.3% accuracy improvement over CoT while using 9.1 fewer tokens per generation and only 1.4 code executions, compared to TIR’s 2.63 code executions. This balance between accuracy and efficiency highlights TATA’s ability to streamline reasoning processes, making it a computationally effective solution for mathematical reasoning tasks.
@@ 4.3 Ablation
Ablation with $\mathcal{H}$
We evaluate the following variants of $\mathcal{H}$: 1. “Random”: This variant uses the same ratio of CoT and TIR as TATA, but the queries are selected randomly. It serves as a strong baseline because it leverages knowledge about the ratio from TATA. 2. “CoT + TIR”: This method includes all CoT and TIR solutions for each query without any data selection. 3. “TATA-”: This variant uses only a single threshold for selection, employing the same strategy for both GSM8K and MATH datasets. More details on these variants can be found in Appendix C.2. From Table 4, we observe that TATA achieves the highest overall accuracy. Although “Random” uses the same ratio of queries relying on either CoT or TIR as TATA, it lags behind by 0.5%, underscoring the importance of the base LLM’s ability to differentiate between when to use CoT or TIR for different queries. Moreover, naively including all CoT and TIR solutions, (i.e. “CoT + TIR”), results in a noticeable decline in performance, despite the larger size of the $\mathcal{D}_{\text{SFT}}$ dataset. This highlights the critical role of effective data selection for CoT and TIR.
Method Random CoT + TIR TATA- TATA
AVG 61.0 58.2 60.2 61.5
Table 4: Ablation Results (in %).
@@ 5.3 Exploring Reinforcement Learning
Recent advancements in reinforcement learning (RL) (OpenAI, 2024; DeepSeek-AI et al., 2025) have demonstrated promising results in enhancing long CoT reasoning. To explore the role of RL in the spontaneous selection between CoT and TIR, we employ Direct Preference Optimization (DPO) to LLMs fine-tuned with our TATA framework (Rafailov et al., 2023) by constructing preference pairs based on the CoT and TIR scores of queries in the new candidate set $\mathcal{D}^{*}={\{(x_{i},y_{i}^{*},z_{i}^{*})\}_{i=1}^{N}}$. Detailed experimental setup and methodologies are provided in Appendix C.5. As shown in Table 6, DPO achieves results comparable to those of TATA. The complete results are provided in Table C.5. This suggests that the original data has already been effectively learned by the base LLM during the SFT stage, and applying additional DPO on the same dataset yields minor improvement. This observation aligns with LIMO (Ye et al., 2025), which argue that the capabilities of pretrained LLMs are latent, with both SFT and RL serving as different methods to elicit these inherent abilities.
@@
6 Related Work
Math Reasoning with CoT and TIR
CoT and TIR are two widely recognized approaches for reasoning with LLMs. CoT offers interpretability and generalizability, while TIR can provide precise calculation results. Previous work on mathematical SFT has primarily focused on either CoT (Yu et al., 2023; Tong et al., 2024; Shao et al., 2024; Yan et al., 2024) or TIR (Yue et al., 2023; Gou et al., 2023; Wang et al., 2023a; Yin et al., 2024), with a few efforts to integrate both (Yue et al., 2023; Beeching et al., 2024; Yang et al., 2024b). For instance, MAmmoTH (Yue et al., 2023) mainly adopts TIR but switches to CoT when code execution fails due to errors or timeouts. However, it relies on separate prompts and manual inference controls to switch between them. Recent work has explored automatic selection between CoT and TIR (Zhao et al., 2023; Yue et al., 2024; Yu et al., 2024), such as using an auxiliary LLM to determine CoT/TIR (Zhao et al., 2023). However, these methods rely on external planners to select CoT/TIR, not by LLMs themselves. In contrast, our work seeks to enable LLMs to spontaneously select the appropriate reasoning strategy without relying on external planners or manual interventions.
Data Selection
Test-Time Scaling
@@ 7 Conclusion
We propose TATA, a novel and effective framework for mathematical reasoning with LLMs that enables models to dynamically align their reasoning strategies, CoT or TIR, with their intrinsic strengths. By incorporating base-LLM-aware data selection during SFT, TATA tailors reasoning strategies to each model, empowering them to select an appropriate paradigm for during inference autonomously. Extensive experiments demonstrate that TATA achieves superior or comparable performance across both in-domain and OOD benchmarks while significantly improving inference efficiency compared to static TIR-based methods. Moreover, our analysis underscores the importance of aptitude-aware data selection in unlocking the potential of LLMs to make autonomous and effective reasoning decisions, paving the way for further advancements in reasoning capabilities of LLMs.
@@ B.4 SFT and Evaluation Setup
We set the maximum input length for all base models to be 4,096. During SFT, we employ the Adam optimizer with a learning rate of $2\times 10^{-5}$ and set batch size to 64, conducting training over three epochs. Unlike Beeching et al. (2024); Yang et al. (2024b), we use the same training prompt for both CoT and TIR. The prompt is provided in Table 9. We utilize 8 A100 GPUs to do SFT, and training a 7B model for 3 epochs with exclusively CoT data spends approximately 12 hours.
@@ Appendix C More Fine-grained Results
C.1 Main Results
The complete results of the main experiments are provided in Table 10. Besides the effectiveness of our proposed framework, the results also reveal distinct behaviors of CoT and TIR across different datasets, highlighting their complementary strengths. CoT excels on in-domain tasks like GSM8K, where step-by-step reasoning is crucial, achieving higher accuracy (e.g., 55.6% for Qwen2.5-0.5B) compared to TIR. However, TIR demonstrates superior performance on symbolic and computation-heavy tasks like MATH, leveraging tool integration to achieve higher accuracy (e.g., 67.5% for Qwen2.5-7B). On out-of-domain datasets, CoT shows strong generalization on structured problems like MAWPS and SVAMP, while TIR performs better on complex tasks like OlympiadBench, where precise computations are essential. This divergence underscores the importance of combining both approaches, as TATA balances CoT’s reasoning strength with TIR’s computational precision, achieving robust performance across diverse datasets.
Model Method Metric In-Domain Out-of-Domain AVG
GSM8K MATH ID AVG MAWPS SVAMP College Olympiad OOD AVG
Qwen2.5-0.5B CoT Acc 55.6 32.5 44.0 86.2 58.9 26.3 6.7 44.5 44.4
Token 305.4 526.8 416.1 195.7 286.2 544.3 768.9 448.8 437.9
# Code 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
TIR Acc 46.9 36.4 41.6 83.5 53.5 26.3 7.7 42.8 42.4
Token 326.2 502.5 414.4 268.6 292.3 492.9 745.8 449.9 438.0
# Code 3.16 2.72 2.94 2.76 2.93 2.91 3.03 2.91 2.92
TATA Acc 52.8 36.6 44.7 85.9 59.4 26.9 8.6 45.2 45.0
Token 309.7 508.7 409.2 217.3 292.9 500.9 743.0 438.5 428.8
# Code 0.19 2.63 1.41 0.52 0.33 2.82 3.06 1.68 1.59
Qwen2.5-1.5B CoT Acc 78.3 48.2 63.2 93.5 82.5 37.9 13.2 56.8 58.9
Token 273.1 488.5 380.8 185.9 233.9 498.9 684.0 400.7 394.0
# Code 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
TIR Acc 70.8 54.3 62.6 91.7 79.8 36.0 19.4 56.7 58.7
Token 324.0 474.5 399.2 275.0 295.4 454.1 689.9 428.6 418.8
# Code 3.1 2.55 2.82 2.83 2.94 2.66 2.96 2.85 2.84
TATA Acc 77.6 53.8 65.7 94.2 80.7 37.0 18.8 57.7 60.4
Token 273.1 480.5 376.8 181.8 235.0 453.3 692.7 390.7 386.1
# Code 0.11 2.37 1.24 0.14 0.21 2.38 2.84 1.39 1.34
Qwen2.5-3B CoT Acc 85.3 53.8 69.6 94.8 85.8 41.5 16.3 59.6 62.9
Token 259.8 469.2 364.5 180.8 230.3 491.3 679.7 395.5 385.2
# Code 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
TIR Acc 80.4 61.4 70.9 89.9 79.8 41.1 25.0 59.0 62.9
Token 312.7 467.0 389.8 283.3 293.8 434.4 676.5 422.0 411.3
# Code 3.04 2.57 2.8 2.94 2.89 2.56 2.81 2.8 2.8
TATA Acc 84.0 61.3 72.6 94.7 85.3 41.6 24.9 61.6 65.3
Token 265.1 468.4 366.8 209.3 237.7 436.4 683.5 391.7 383.4
# Code 0.07 2.44 1.25 0.55 0.26 2.49 2.8 1.52 1.43
Qwen2.5-7B CoT Acc 88.7 58.6 73.6 96.1 88.1 42.7 23.0 62.5 66.2
Token 255.8 464.5 360.2 177.7 217.9 480.0 673.5 387.3 378.2
# Code 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
TIR Acc 86.8 67.5 77.2 92.1 84.3 44.2 31.9 63.1 67.8
Token 306.4 452.3 379.4 261.6 272.0 430.6 636.3 400.1 393.2
# Code 3.01 2.43 2.72 2.58 2.54 2.52 2.68 2.58 2.63
TATA Acc 89.5 66.8 78.2 94.2 86.2 43.4 31.1 63.7 68.5
Token 258.1 452.8 355.5 195.3 222.0 433.9 652.4 375.9 369.1
# Code 0.18 2.27 1.23 0.5 0.29 2.49 2.66 1.48 1.4
LLaMA-3-8B CoT Acc 84.7 46.5 65.6 91.6 81.6 30.2 13.3 54.2 58.0
Token 246.4 471.0 358.7 173.3 236.8 511.7 676.7 399.6 386.0
# Code 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
TIR Acc 81.7 56.2 69.0 87.8 77.8 30.5 21.9 54.5 59.3
Token 299.0 457.5 378.2 240.9 269.1 437.9 650.8 399.7 392.5
# Code 2.96 2.51 2.74 2.42 2.64 2.69 2.76 2.63 2.66
TATA Acc 84.0 55.1 69.6 91.8 82.7 34.2 21.5 57.6 61.5
Token 248.2 461.1 354.6 191.1 222.5 449.5 657.7 380.2 371.7
# Code 0.12 2.33 1.23 0.27 0.21 2.39 2.6 1.37 1.32
Qwen2.5Math-1.5B CoT Acc 83.1 56.6 69.8 92.9 84.1 44.3 19.3 60.2 63.4
Token 267.1 476.5 371.8 181.7 238.3 485.7 681.7 396.8 388.5
# Code 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
TIR Acc 78.8 64.8 71.8 92.3 83.3 42.4 27.4 61.4 64.8
Token 332.2 516.6 424.4 324.5 335.6 479.0 772.5 477.9 460.1
# Code 3.28 2.8 3.04 3.62 3.55 2.99 3.14 3.32 3.23
TATA Acc 83.2 62.8 73.0 94.0 85.6 43.9 26.8 62.6 66.0
Token 266.2 508.9 387.5 199.4 247.1 467.1 743.6 414.3 405.4
# Code 0.26 1.68 0.97 0.32 0.42 1.44 2.38 1.14 1.08
Qwen2.5Math-7B CoT Acc 91.0 61.5 76.2 94.8 87.9 45.7 23.9 63.1 67.5
Token 254.7 470.6 362.6 177.0 223.5 484.1 669.2 388.5 379.9
# Code 0.0 0.0 0.0 0.0 0.0 0.0 0.01 0.0 0.0
TIR Acc 88.9 73.6 81.2 95.4 89.4 47.1 35.3 66.8 71.6
Token 311.8 490.9 401.4 261.2 272.2 456.8 713.7 426.0 417.8
# Code 3.04 2.56 2.8 2.58 2.51 2.65 2.75 2.62 2.68
TATA Acc 89.8 73.0 81.4 95.2 88.1 48.3 35.9 66.9 71.7
Token 264.7 487.2 376.0 193.7 229.7 476.9 710.6 402.7 393.8
# Code 0.25 2.14 1.2 0.33 0.24 2.02 2.59 1.3 1.26
Table 10: Detailed results of our TATA framework. The best accuracies within each group are shown in bold. The three metrics, “Acc”, “Token”, and “# Code” represent the average accuracy (%), total tokens per generation, and number of code executions. “ID AVG”, “OOD AVG”, and “AVG” denote the averages of these metrics across in-domain, out-of-domain, and all six benchmarks. “CoT”, “TIR”, and “TATA” indicate fine-tuning exclusively on CoT data, TIR data, and using our TATA framework, respectively, with the corresponding base LLM.
C.2 Ablation Study
@@
C.3 Analysis of CoT scores and TIR scores

### v3

@@ 5.1 Experimental Setup
We include the following methods as our baselines: 1) Hybrid (Yue et al., 2023): Primarily uses TIR but falls back to CoT upon code execution errors or timeouts (Figure 1 (b)). 2) Ensemble (Zhao et al., 2023): Post-selects between TIR and CoT outputs using an additional LLM (Figure 1 (a)). In our implementation, we use the same 8-shot prompt as Zhao et al. (2023) with the base LLM as the selector for consistency. 3) GPT-Select: Uses GPT-4o during data selection to choose CoT or TIR per query, testing whether a strong external LLM can effectively select reasoning paradigms regardless of the base LLM’s aptitude.
@@ 5.2 Main Results
Effectiveness of TATA
Results presented in Table 1 demonstrate the effectiveness of our proposed TATA framework. Across various base models, model sizes, and benchmarks, TATA consistently achieves competitive or superior performance compared to all the other baselines, highlighting its ability to leverage the complementary advantages of both methods. Additionally, TATA achieves significantly better performance than the “GPT-Select” baseline. While “GPT-Select” leverages a much stronger LLM to select between CoT and TIR for different queries, it demonstrates that this approach may not be suitable for all base LLMs. This highlights the critical importance of base-LLM-aware selection in optimizing performance.
@@
The results in Table 2 demonstrate that our TATA not only improves accuracy but also enhances inference efficiency compared to standalone CoT and TIR methods. Across all model sizes, TATA achieves higher accuracy while maintaining lower token usage and fewer code executions than TIR, and it significantly reduces computational overhead compared to TIR without sacrificing the benefits of tool integration. For instance, with Qwen2.5-7B, TATA achieves a 2.3% accuracy improvement over CoT while using 9.1 fewer tokens per generation and only 1.4 code executions, compared to TIR’s 2.63 code executions. This balance between accuracy and efficiency highlights TATA’s ability to streamline reasoning processes, making it a computationally effective solution for mathematical reasoning tasks. The “hybrid” and “ensemble” approaches incur even higher inference costs compared to our proposed TATA. Specifically, "hybrid" requires decoding via TIR and selectively switching to CoT execution for specific cases; “ensemble” generates both CoT and TIR outputs during testing and incurs additional costs for selection between the two.
@@ 5.3 Ablation
Anchor set & Others
Table 3 includes results for several other ablation studies: 1) “CoT + TIR”: This method includes all CoT and TIR solutions for each query without any data selection. 2) Anchor set construction with random sampling ("TATA - random 100"): Replacing k-means clustering with random selection while keeping the anchor set size constant. 3) Larger anchor set size ("TATA - A=200"): Increasing the anchor set size to 200. From Table 3, we observe that TATA achieves the highest overall accuracy. Naively including all CoT and TIR solutions (i.e., “CoT + TIR”) results in a noticeable decline in performance, despite the larger size of the $\mathcal{D}_{\text{SFT}}$ dataset. Random anchor set selection ("TATA - random 100") critically degrades performance, highlighting the importance of a representative anchor set over size alone. Increasing the anchor set size shows diminishing returns, indicating that $A=100$ is enough for model evaluation in our SFT data curation.
@@ 6.3 Exploring Reinforcement Learning
Recent advancements in RL (OpenAI, 2024; DeepSeek-AI et al., 2025) have demonstrated promising results in enhancing long CoT reasoning. To explore the role of RL in the spontaneous selection between CoT and TIR, we employ Direct Preference Optimization (DPO) to LLMs fine-tuned with our TATA framework (Rafailov et al., 2023) by constructing preference pairs based on the CoT and TIR scores of queries in the new candidate set $\mathcal{D}^{*}={\{(x_{i},y_{i}^{*},z_{i}^{*})\}_{i=1}^{N}}$. Detailed experimental setup and methodologies are provided in Appendix C.4. As shown in Table 6, DPO achieves results comparable to those of TATA. The complete results are provided in Table C.4. This suggests that the original data has already been effectively learned by the base LLM during the SFT stage, and applying additional DPO on the same dataset yields minor improvement. This observation aligns with LIMO (Ye et al., 2025), which argues that the capabilities of pretrained LLMs are latent, with both SFT and RL serving as different methods to elicit these inherent abilities.
@@ 7 Conclusion
We propose TATA, a novel and effective framework for mathematical reasoning with LLMs that enables models to dynamically align their reasoning strategies, CoT or TIR, with their intrinsic strengths. By incorporating base-LLM-aware data selection during SFT, TATA tailors reasoning strategies to each model, empowering them to select an appropriate paradigm for inference autonomously. Extensive experiments demonstrate that TATA achieves superior or comparable performance across both in-domain and OOD benchmarks while significantly improving inference efficiency compared to method based on TIR alone. Moreover, our analysis underscores the importance of aptitude-aware data selection in unlocking the potential of LLMs to make autonomous and effective reasoning decisions, paving the way for further advancements in reasoning capabilities of LLMs.
@@ B.4 SFT and Evaluation Setup
We set the maximum input length for all base models to be 4,096. During SFT, we employ the Adam optimizer with a learning rate of $2\times 10^{-5}$ and set batch size to 64, conducting training over three epochs. Unlike Beeching et al. (2024); Yang et al. (2024b), we use the same training prompt for both CoT and TIR. The prompt is provided in Table 9.
@@ Appendix C More Fine-grained Results
C.1 Ablation Study
@@
C.2 Analysis of CoT scores and TIR scores